OpenStudy (volleyballlover55):

What is the rational number equivalent to 3 point 12 with a bar over 12? 3 and 4 over 33 3 and 8 over 33 3 and 10 over 39 3 and 5 over 39

OpenStudy (volleyballlover55):

@welshfella

OpenStudy (welshfella):

i'm trying to remember how do do these..... 100 * 0.1212.. = 12 .1212.... 1 * 0.1212... = 0 . 1212... subtract 99 * 0.1212... = 12 so its 12/99

OpenStudy (welshfella):

if you enter 12 / 99 into your calculator you'll get 0.1212....

OpenStudy (volleyballlover55):

ok.. but whats the answer?

OpenStudy (welshfella):

well you need to put 3 in the front of course then simplify 12/99

OpenStudy (welshfella):

- divide top and bottom by 3

OpenStudy (welshfella):

12 /3 and 99 / 3

OpenStudy (volleyballlover55):

ok 4/33

OpenStudy (welshfella):

right so the answer is ?

OpenStudy (volleyballlover55):

A!

OpenStudy (welshfella):

right